3.) In a political rally, an opinion poll was conducted among the present voters if they approve the platform of the candidate. 201 said they do approve. If there were 500 individuals in the sample, what is p? what is q?


4.) A certain barangay embarked on a Quit Smoking program among the constituents. A survey was made among a sample of 200 smoking individuals of the village asking who would join the program. 118 said they would join the program. What are the values of p and q?


5.) A total of 200 Grade 8 students who have access to internet services were asked if they play games online before they attend classes. 95 students responded Yes. What is the population proportion of students who play video games before attending classes? What percent do not play video games before attending classes?


1
Expert's answer
2021-06-07T04:45:09-0400

3)

"x=201 \\

n=500 \\

p= \frac{x}{n}=\frac{201}{500}=0.402 \\

q = 1-p = 1-0.402=0.598"

4)

"n=200 \\

x=118 \\

p=\frac{118}{200}=0.59 \\

q=1-0.59=0.41"

5)

"n=200 \\

x=95 \\

p=\frac{95}{200}=0.475 \\

q=1-0.475=0.525"

or 52.5 %